carmelle:

Find the equation of a line that is perpendicular to line g that contains (P, Q). A. 3x − y = 3P − Q B. 3x + y = Q − 3P C. x − y = P − Q D. x + y = Q − P

carmelle:

Vocaloid:

Determine the slope of g (visually, you can tell the rise/run) Perpendicular —> take the opposite reciprocal of the slope —> -1/slope from the previous step set up y - y1 = m(x-x1) using point P Q as your x1, y1 Plug in the m from the opposite reciprocal step Re-arrange the equation if needed, to get one of the answer choices

carmelle:

alright, so here's what i've got (ive been trying to do this but im kinda confused on where i went wrong..?) slope of line g: -1/3 perpendicular line slope: 3 (?) y - Q = 3(x - P) y - Q = 3x - 3P -3x + y - Q = 3P -3x + y = 3P + Q Steps: 1. Given, in point-slope form 2. Distribute 3. -3x on both sides 4. +Q on both sides what do I do next, that's not one of the options?? did I mess up?

Vocaloid:

nothing's incorrect, just need to play around a bit more w/ the re-arrangement step from here: y - Q = 3x - 3P subtracting y from both sides adding 3P to both sides 3x - y = 3P - Q (choice A)

carmelle:

ohhh ok, tysm!! ^